Two closed vessels of equal volume 8.21L are connected by narrow tube of negligible volume with open valve. The LHS container is found to contain 3 mole of CO2 &2 mole of He at 400k , what is partial pressure of He in RHS container at 500k.

Grade:11

1 Answers

Dreamer artists
5 years ago
Total no of moles = 2+3 = 5 
Total pressure p = nRT/V
Partial pressure of He = nHe/not × p
= 2/5* 5*0.0821*500/8.21 = 10 atm
Thank you 
Hope it helps.